OpenStudy (anonymous):

Given: mÐ1 = mÐ3 Prove: mÐAFC = mÐDFB

OpenStudy (anonymous):

mÐ1 = mÐ3 a. Given mÐ1 + mÐ2 = mÐ3 + mÐ2 b. Addition Property of Equality mÐ1 + mÐ2 = mÐAFC mÐ3 + mÐ2 = mÐDFB c. __________ ÐAFC = ÐDFB d. Substitution Property

OpenStudy (anonymous):

Select one: a. Addition Property of Equality b. Angle Addition Postulate c. Multiplication Property of Equality d. Division Property of Equality

OpenStudy (anonymous):

@dan815 : can you help me please ?